All please

All please
media%2F929%2F929562d4-afcc-4487-b4d0-02
Answer #1

(a) the missing side is

x2 = 42 + 52

x = V41

so,

sin heta = 4V41 / 41

(B) is correct option

-----------------------------------------------------------------------------------------------------------

(b) the angle will be

alpha = sin-1 (0.18293545)

alpha = 10.54

(c) alpha = tan-1 (0.91203476)

alpha = 42.36

----------------------------------------------------------------------------------------

130.6 degree

multiply by pi / 180

we get 2.279 radians